A high frequency signal source with large output resistance is to be connected to the
input of an op-amp inverting amplifier. Which characteristics of the op-amp will be
significant in handling such signal? Justify your answer.
1
Expert's answer
2019-02-20T13:50:36-0500

The amplifiers are created to amplify signals within the given bandwidth.

At high frequencies the capacitor' reactance decreases according to


"X_C=\frac{1}{2\pi fC},"


so it may seem that we should use higher "R_{in}". But the signal source already has high resistance, so it will cover low value of "R_{in}". Moreover, as the frequency increases, the closed-loop gain begins to drop, so to avoid it according to its definition,

"G_{CL} \approx -\frac{R_f}{R_{in}},"


take low "R_{in}".

What about R feedback? It shouldn't be too high because at very high values (megaohms) the bandwidth drops off thus limiting the frequency response and making noise go up.